OpenStudy (anonymous):

in each game of sic bo, three regular six sided dice are throw once. (a) what is the probability that exactly two dice show six? (b) what is the probability that exactly two of the dice show the same number? (c) what is the probability of rolling three different numbers on the dice

OpenStudy (dumbcow):

there are 6^3 total possible outcomes there are 6*3 outcomes where 2 dice show six. probability = 18/216

OpenStudy (dumbcow):

what is sic bo ??

OpenStudy (anonymous):

i have no idea lol i think the text book made up a game

OpenStudy (dumbcow):

haha

OpenStudy (dumbcow):

i think i may have been wrong 1) = 15/216 2) = 90/216 3) =120/216

OpenStudy (anonymous):

thanks heaps :)